OpenStudy (anonymous):

if the cost price of the article is 30% of the selling price. what percentage of cost price is the selling price. A- 230% B- 70% C- 330% D- 130%

OpenStudy (wolf1728):

cost = .30 * selling price 1/.30 = 3.33 So, 3.33 * cost = selling price Answer is 333% which is almost answer C

OpenStudy (anonymous):

hmm why do you divide 1 by .33

OpenStudy (perl):

when you see the % symbol, immediately divide by 100

OpenStudy (perl):

Let c = cost price, s = selling price c = 30% * s = 30/100 * s = 0.3 * s now you want to solve Translate the english into algebra "what percentage of the cost price is the sales price" the percentage is unknown, call it x , x % * c = s x/100 * c = s solve for x 100 s = x *c 100 s / c substitute c = .30 * s 100 s / (.30 s ) = x cancel s 100 / .30 = x x = 333.33 %

OpenStudy (anonymous):

wolf1728 thanks yes its right